the steps to find the answer of 3 over 5 =-12
3/5x= -12

Answer by Theo(13342)

(3/5) * x = - 12

multiply both sides of the equation by 5 to get:

3*x = -12*5

simplify to get:

3*x = -60

divide both sides of the equation by 3 to get:

x = -20

(3/5) * -20 = -60/5 = -12 so you're good to go.